OpenStudy (anonymous):

The sum of three consecutive number is eighty - four. What is the smallest of the three numbers?

OpenStudy (p0sitr0n):

n+(n+1)+(n+2)=84 3n+3=84 n=27 n+1 = 28 n+2 = 29 27+28+29=84